OpenStudy (anonymous):

What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7 fan for best answer can someone really help me im going over all the test i failed and i failed this test twice both got a 70% so some one plz help me

OpenStudy (anonymous):

So what you are trying to do is figure out what make a + b + c equal to 0

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I'm not done explaining

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

So we want to simplify for a + (–7) + 3 = 0 so we get this\[a +-4 = 0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is also written as a - 4 = 0 wat we want to do now is add four to both sides so that you get a alone

OpenStudy (anonymous):

@superman36

OpenStudy (anonymous):

the answer is a = 4

OpenStudy (anonymous):

so 4 + (-7) + 3 = 0

OpenStudy (anonymous):

@mariah232323

OpenStudy (anonymous):

Medal please?